# Lines of Thought in Large Language Models

### v2

@@ 5 Conclusion
That’s not to say that reduced complexity representations are necessarily useful in practice. Individual trajectory variability persists, and is essential to accurately predict next tokens and continue textual inputs. Thus, it is not immediately apparent to us whether the low-dimensionality structures identified could lead to avenues for compressing or ablating transformers, although it might.
Yet, patterns are explanatory. Our concern here has been primarily to discover some of the mechanisms implicitly encoded in the weights of trained language models. Further investigations could extend this methodology to more thoroughly identify and characterize the dynamics of tokens.
@@ A.2 Langevin equation derivation
with $\Lambda(t,\tau)=\mathbf{\Sigma}(t+\tau)\mathbf{\Sigma}^{-1}(t)$, and assuming now that $t,\tau$ are variables in $\mathbb{R}$, as $\tau$ goes to 0 we can approximate:
@@ A.3 Numerical integration
where $\alpha=(t-t_{1})/(t_{2}-t_{1})$ and with $\ln_{\mathrm{M}},\exp_{\mathrm{M}}$ denoting the matrix logarithm and exponential, respectively.181818 $\exp_{\mathrm{M}}({\bm{A}})=\sum{\bm{A}}^{k}/k!$ and $\ln_{M}$ is the inverse function: $\ln_{\mathrm{M}}\left[\exp_{\mathrm{M}}({\bm{A}})\right]={\bm{I}}$. This also yields the derivative $\dot{{\bm{U}}}(t)=\left[{\bm{U}}\ln_{M}{{\bm{R}}}\right]/(t_{2}-t_{1})$. Indeed:
@@ B.1 Trajectory clustering
In Fig. 6, we show evidence of trajectory clustering in the latent space. In particular, all pilot tokens get transported away from the origin (or their starting point) by a comparable amount, resulting in narrow distributions along the first singular direction. Another signature of clustering is the fact that token positions at different times form distinct clusters, as showed by low-dimensional t-SNE representation (van der Maaten & Hinton, 2008).
Figure 6: (Left) Distributions along the first singular vector at different times. (Right) Low-dimensional (t-SNE) visualization of the clustering of tokens, notably across different times. Same color legend.
@@ B.2 Trajectory extrapolation
In Fig. 7, we provide a schematic to explain the reasoning behind Eq. 1. If the cluster rotated and stretched like a solid, the position of a point ${\bm{x}}^{\prime}$ at $t^{\prime}$ could be inferred exactly from it position ${\bm{x}}$ at $t$, using the formula outlined. However, unsurprisingly, the token ensemble does not maintain its topology and the points move around the clusters, requiring the stochastic term ${\mathbf{w}}$ injected in Eq. 1.
Figure 7: Extrapolation between $t$ and $t^{\prime}$. The extrapolated location ${\bm{x}}^{\prime}$ corresponds to the rotated and stretched position of ${\bm{x}}$. Given that $\vec{u}={\bm{R}}\vec{e}$, $\vec{u}^{\prime}={\bm{R}}^{\prime}\vec{e}$ and ${\bm{R}}^{-1}={\bm{R}}^{\top}$, we have $\vec{e}={\bm{R}}^{\top}\vec{u}={\bm{R}}^{\prime\top}\vec{u}^{\prime}$ and thus $\vec{u}^{\prime}={\bm{R}}^{\prime}{\bm{R}}^{\top}\vec{u}$.
B.3 Noise statistics
Fig. 8 provides additional details pertaining to the distribution of residuals $\delta_{x}$. Since they are many dimensions and time points, it gives only representative snapshots. It intends to substantiate the results that:
  • the $\delta{\bm{x}}$ are Gaussian (Fig. 8A);
  • the variance is exponential in $(t+\tau)$, with no dependency on $t$ (Fig. 8B);
  • all components $\delta x_{i}$ of $\delta{\bm{x}}$ have the same distribution (Fig. 8C), i.e., isotropy;
  • there are no spatial cross-correlations, i.e. $\langle\delta x_{i}\delta x_{j}\rangle=\delta_{ij}$ (Dirac function) (Fig. 8D).
Figure 8: Statistics of $\delta\mathbf{x}$. (A) Empirical PDF of $\delta x_{42}(10,t+\tau)$, with $t+\tau=12,14,16$. The curves appear Gaussian. (B) Variance of $\delta x_{i}$ for $i=1\dots 8$, for $t=4,8,12,16$ and $t+\tau>t$. (C) Empirical PDF of $\delta x_{i}(12,14)$ for $i=1\dots 1024$. The curves are similar for almost all coordinates. (D) Cross correlations of $\delta x_{i}$ and $\delta x_{j}$.
B.4 Details on noise aggregated statistics (Fig. 4)
@@ B.3 Noise statistics
Figure 9: Schematic to explain the noise figures such as Fig. 4. Each square represents a summary statistics. Specifically, the square at $(t,t+\tau)$ represents the distribution of $\{\delta{\bm{x}}^{(k)}(t,t+\tau)\}_{k}$, with $k$ indexing individual tokens. The $\delta{\bm{x}}$ along each coordinate $i$ form a distribution, from which one can extract the corresponding $\mu_{i},\sigma_{i},\kappa_{i}$ (mean, variance, kurtosis). These 1D moments are then averaged along all coordinates $i$ ($\langle\mu_{i}\rangle_{i}$, etc.), forming the value displayed in the square.
B.5 Null testing
Fig. 10 shows the trajectories of language vs gibberish, as well as the linear separability of the two ensemble. It also shows trajectories for an untrained GPT-2 shell, and a model with only the last 12 layers reinitialized.
Figure 10: (Top-left) Trajectories of non-language (red) vs language (black), plotted in the same axes (10-token pseudo-sentences). (Top-right) Accuracy of linear separability between language and non-language for each layer. Obtained by training a Perceptron (train/test: 0.7/0.3; 14000 trajectories). (Bottom-left) Trajectories in the untrained GPT-2 model. They are transported in straight lines. (Bottom-right) Trajectories in the mixed model. After being transported by trained layers 1-12, the trajectories stop. Layers 13-24 with random weights do not transport tokens any further.
B.6 Results with other models
@@
Figure 11: GPT-2 untrained. The averaged excess kurtoses $\langle|\kappa|\rangle$ fall in the 11.5 range, indicating strong non-gaussianity. The variance does not scale solely with $t+\tau$.
@@
Figure 13: Mistral 7B v0.1. The last layer (32) appears to have an anomalously large variance.
@@
Figure 14: Llama 3.2 1B. This small model all present an out-of-distribution last layer.
@@
Figure 15: Llama 3.2 3B. The last layer anomaly is also present.

### v3

@@ 5 Conclusion
Patterns are explanatory. Our concern here has been primarily to discover some of the mechanisms implicitly encoded in the weights of trained language models. Yet, there are also concrete and potentially practical implications to our findings.
For interpretability, finding low-dimensional structures is consequential. It is one of the most efficient ways to break down the inherent complexity of large models into more elementary constituents. Our dynamical system approach reveals a surprising dimensionality reduction of token embeddings. It suggests, notably, that the true “meaning” of embeddings is contained within individual variability (possibly orthogonal to the average pathway collectively followed by all LoTs). This is also merely a first-order approximation, which could be extended to more complete and precise equations, where the “noise term” becomes smaller and smaller. Eventually, we anticipate the possibility for hybrid architectures where the deterministic part of trajectories is delegated to a small system of equations, while the variable part, where meaning is encoded, is handled by a neural network; potentially with many fewer weights.
Our theoretical model in Eqs. 3 and 4 not only reveals low-dimensionality, but also extends token trajectories to continuous paths. In the past, the Neural Ordinary Differential Equation paradigm by Chen et al. (2019) showed that converting a discrete neural network into a continuous dynamical system had many advantages. Notably, it offers opportunities for compression and stability, while pointing towards efficient hybrid architectures. Our paper demonstrates that transformers can also been seen through the lens of dynamical systems, with a similar continuous extension as seen in neural ODEs.
Finally, the new methodology that we introduced is portable and widely applicable. Incidentally, it can also serve as a diagnostic method to highlight intrinsic differences between transformer layers. Fig. A8 to Fig. A10, for example, show significant deviations in the last layer (and to a lesser extent in the early ones). This suggests that these layers achieve a different kind of processing than intermediate layers, possibly following fine-tuning and/or re-alignment. It’s not immediately obvious to us how these “anomalies” could be detected through a different approach.
@@ A.2 Langevin equation derivation
with $\mathbf{\Lambda}(t,\tau)=\mathbf{\Sigma}(t+\tau)\mathbf{\Sigma}^{-1}(t)$, and assuming now that $t,\tau$ are variables in $\mathbb{R}$, as $\tau$ goes to 0 we can approximate:
@@ A.3 Numerical integration
where $\alpha=(t-t_{1})/(t_{2}-t_{1})$ and with $\ln_{\mathrm{M}},\exp_{\mathrm{M}}$ denoting the matrix logarithm and exponential, respectively.171717 $\exp_{\mathrm{M}}({\bm{A}})=\sum{\bm{A}}^{k}/k!$ and $\ln_{M}$ is the inverse function: $\ln_{\mathrm{M}}\left[\exp_{\mathrm{M}}({\bm{A}})\right]={\bm{I}}$. This also yields the derivative $\dot{{\bm{U}}}(t)=\left[{\bm{U}}\ln_{M}{{\bm{R}}}\right]/(t_{2}-t_{1})$. Indeed:
@@ B.1 Trajectory clustering
In Fig. A1, we show evidence of trajectory clustering in the latent space. In particular, all pilot tokens get transported away from the origin (or their starting point) by a comparable amount, resulting in narrow distributions along the first singular direction. Another signature of clustering is the fact that token positions at different times form distinct clusters, as showed by low-dimensional t-SNE representation (van der Maaten & Hinton, 2008).
Figure A1: (Left) Distributions along the first singular vector at different times. (Right) Low-dimensional (t-SNE) visualization of the clustering of tokens, notably across different times. Same color legend.
@@ B.2 Trajectory extrapolation
In Fig. A2, we provide a schematic to explain the reasoning behind Eq. 1. If the cluster rotated and stretched like a solid, the position of a point ${\bm{x}}^{\prime}$ at $t^{\prime}$ could be inferred exactly from it position ${\bm{x}}$ at $t$, using the formula outlined. However, unsurprisingly, the token ensemble does not maintain its topology and the points move around the clusters, requiring the stochastic term ${\mathbf{w}}$ injected in Eq. 1.
Figure A2: Extrapolation between $t$ and $t^{\prime}$. The extrapolated location ${\bm{x}}^{\prime}$ corresponds to the rotated and stretched position of ${\bm{x}}$. Given that $\vec{u}={\bm{R}}\vec{e}$, $\vec{u}^{\prime}={\bm{R}}^{\prime}\vec{e}$ and ${\bm{R}}^{-1}={\bm{R}}^{\top}$, we have $\vec{e}={\bm{R}}^{\top}\vec{u}={\bm{R}}^{\prime\top}\vec{u}^{\prime}$ and thus $\vec{u}^{\prime}={\bm{R}}^{\prime}{\bm{R}}^{\top}\vec{u}$.
B.3 Separability of true and extrapolated positions
To characterize the similarity between the ensemble of true positions at $t+\tau$, ${\bm{x}}(t+\tau)$, from the positions extrapolated from $t$, $\tilde{{\bm{x}}}(t,\tau)={\bm{R}}(t+\tau)\mathbf{\Lambda}(t,\tau){\bm{R}}(t)^{\top}{\bm{x}}(t)$, we evaluate how much the two sets can be separated with a linear classifier. We train a Support Vector Machine Model with a linear kernel for each set of extrapolations $\{\tilde{{\bm{x}}}(t,\tau)\}$ (70/30 train/test). We then apply the classifier to predict whether points in the test set are true or extrapolated. In Table 1, we report results for the panels corresponding to Fig. 3. The accuracy of the classifier lies in the 50%-60% range, barely above random guessing (50%).
Table 1: Accuracy of linear classifier to separate ${\bm{x}}(t+\tau)$ and $\tilde{{\bm{x}}}(t,\tau)$, in percent.
$t=12$ $t=14$ $t=16$ $t=18$
$t+\tau=13$ 48
$t+\tau=15$ 51 49
$t+\tau=17$ 56 53 47
$t+\tau=19$ 54 56 55 46
$t+\tau=21$ 56 60 61 55
B.4 Noise statistics
Fig. A3 provides additional details pertaining to the distribution of residuals $\delta_{x}$. Since they are many dimensions and time points, it gives only representative snapshots. It intends to substantiate the results that:
  • the $\delta{\bm{x}}$ are Gaussian (Fig. A3A);
  • the variance is exponential in $(t+\tau)$, with no dependency on $t$ (Fig. A3B);
  • all components $\delta x_{i}$ of $\delta{\bm{x}}$ have the same distribution (Fig. A3C), i.e., isotropy;
  • there are no spatial cross-correlations, i.e. $\langle\delta x_{i}\delta x_{j}\rangle=\delta_{ij}$ (Dirac function) (Fig. A3D).
Figure A3: Statistics of $\delta\mathbf{x}$. (A) Empirical PDF of $\delta x_{42}(10,t+\tau)$, with $t+\tau=12,14,16$. The curves appear Gaussian. (B) Variance of $\delta x_{i}$ for $i=1\dots 8$, for $t=4,8,12,16$ and $t+\tau>t$. (C) Empirical PDF of $\delta x_{i}(12,14)$ for $i=1\dots 1024$. The curves are similar for almost all coordinates. (D) Cross correlations of $\delta x_{i}$ and $\delta x_{j}$.
B.5 Details on noise aggregated statistics (Fig. 4)
@@ B.4 Noise statistics
Figure A4: Schematic to explain the noise figures such as Fig. 4. Each square represents a summary statistics. Specifically, the square at $(t,t+\tau)$ represents the distribution of $\{\delta{\bm{x}}^{(k)}(t,t+\tau)\}_{k}$, with $k$ indexing individual tokens. The $\delta{\bm{x}}$ along each coordinate $i$ form a distribution, from which one can extract the corresponding $\mu_{i},\sigma_{i},\kappa_{i}$ (mean, variance, kurtosis). These 1D moments are then averaged along all coordinates $i$ ($\langle\mu_{i}\rangle_{i}$, etc.), forming the value displayed in the square.
B.6 Null testing
Fig. A5 shows the trajectories of language vs gibberish, as well as the linear separability of the two ensemble. It also shows trajectories for an untrained GPT-2 shell, and a model with only the last 12 layers reinitialized.
Figure A5: (Top-left) Trajectories of non-language (red) vs language (black), plotted in the same axes (10-token pseudo-sentences). (Top-right) Accuracy of linear separability between language and non-language for each layer. Obtained by training a Perceptron (train/test: 0.7/0.3; 14000 trajectories). (Bottom-left) Trajectories in the untrained GPT-2 model. They are transported in straight lines. (Bottom-right) Trajectories in the mixed model. After being transported by trained layers 1-12, the trajectories stop. Layers 13-24 with random weights do not transport tokens any further.
B.7 Results with other models
@@
Figure A6: GPT-2 untrained. The averaged excess kurtoses $\langle|\kappa|\rangle$ fall in the 11.5 range, indicating strong non-gaussianity. The variance does not scale solely with $t+\tau$.
@@
Figure A8: Mistral 7B v0.1. The last layer (32) appears to have an anomalously large variance.
@@
Figure A9: Llama 3.2 1B. This small model all present an out-of-distribution last layer.
@@
Figure A10: Llama 3.2 3B. The last layer anomaly is also present.